OpenStudy (anonymous):

0.4x + 3y = 2.6 x - 2y = 4.6

OpenStudy (hba):

@Chloeee_Harris Got it ?

OpenStudy (anonymous):

not really :/

OpenStudy (australopithecus):

Use system of equations: For example: 3x + 2 + y = 3 y + 5 + x = 1 1. solve for a variable in one of your equations (I'm using y + 5 + x = 1) 4 + x = y 2. sub into the other equation (I'm subbing it into this 3x + 2 + y = 3) 3x + 2 + (4 + x) = 3 3. Solve for x and sub what you find x to be back into one of the original two equations and then solve for y Note: what you are doing here is determining where the two equations intersect, thus if both of the graphs are parallel to each other system of equations will have no solutions In this example we are finding the red dot in the graph found at this link (the intersection of these two equations like I said) http://openstudy.com/study#/updates/50d71652e4b0d6c1d5420b1d

OpenStudy (australopithecus):

If these graphs had the same slope they would not intersect thus they would have no solutions, for example take the equation of the line: y = mx + b where, m = slope b = y-intersect EXAMPLE 1: just say both graphs have a slope of 2, and one had an intersect of 5 and another an intersect of 7 y = 2x + 5 y = 2x + 7 see from both of their graphs you can see they have no intersect thus there is no solution to the system of equations https://www.wolframalpha.com/input/?i=graph++y+%3D+2x+%2B+5+and+y+%3D+2x+%2B+7 EXAMPLE 2: Just say you have two of the same equations y = 2x + 7 y = 2x + 7 If you apply system of equations to them you will find that they are equal at all points (because they are the same equation they will just overlap if you graph them) ------------------------------------------------------------------------ These are all the possible cases for equation of the line, excluding if you are using equations that intersect multiple times, but those are straight forward and the first example pretty much covers solving them to. If you have any questions please dont hesitate to ask

OpenStudy (anonymous):

?

OpenStudy (australopithecus):

Do you not follow?

OpenStudy (australopithecus):

take a time and read what I wrote

OpenStudy (anonymous):

this is simultaneous equations?

OpenStudy (anonymous):

im not solving graphically

OpenStudy (australopithecus):

yes you would use system of equations to solve this problem, I know you aren't but it is good to understand what you are doing with system of equations, so you dont get tripped up on a test also it has real world applications.

OpenStudy (anonymous):

im so confuzed, im supposed to be A grade mathematics however i can barely understand this :(

OpenStudy (australopithecus):

just read it carefully and think about it

OpenStudy (australopithecus):

in the 3 steps I show you how to solve a problem just like yours

OpenStudy (australopithecus):

also note that both equations have different slopes so they intersect

OpenStudy (australopithecus):

If you have problems with the algebra feel free to post your steps and I can check them for you or show you the rule to apply

OpenStudy (australopithecus):

Remember Slope is the rate of change of an equation, so as you type larger x numbers into the equation the equation will change by the amount dictated as the slope. Thus it is reasonable to assume that when two equations have different slopes that one will increase faster than the other and thus will have to intersect it eventually (when it comes to lines or equations with the formula y=mx + b)

OpenStudy (australopithecus):

Fix: *Remember Slope is the rate of change of an equation, so as you type larger x numbers into the equation the equation y will change by the amount dictated by the slope. In this case x is your input variable and y is your output variable so you sub a number into x and you get a number y

OpenStudy (australopithecus):

@Chloeee_Harris For the first post I made I meant to post this link, not the link to openstudy please check it out, it is the graph of 3x + 2 + y = 3 and y + 5 + x = 1 https://www.wolframalpha.com/input/?i=graph+3x+%2B+2+%2B+y+%3D+3+and+y+%2B+5+%2B+x+%3D+1

OpenStudy (australopithecus):

you can see the red dot that I was referring to in the last post

OpenStudy (australopithecus):

Sorry about that

OpenStudy (anonymous):